OpenStudy (anonymous):

Which equation demonstrates the opposite of a sum property? A. (22 + 10) + 3 = (10 + 22) + 3 B. –5 + 13 = 13 + (–5) C. (12 + (–4)) + 14 = 12 + ((–4) + 14) D. –(–6 + 17) = 6 + (–17)

OpenStudy (landon34):

I got a cool website you can check out for help. here is the URL

OpenStudy (anonymous):

Can anyone help :c

OpenStudy (landon34):

http://www.geteasysolution.com/

OpenStudy (welshfella):

This must be a US thing . I've never heard of it.

OpenStudy (anonymous):

:/

OpenStudy (landon34):

did you even see my answer? not trying to be mean, if it sounds mean.

OpenStudy (welshfella):

http://www.ask.com/math/property-opposite-sum-59c9086dd930e208 - this is a site you can look at

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

But, I'm gonna need someone to actually explain it to me...

OpenStudy (anonymous):

But thanks for your help anyway :P

OpenStudy (welshfella):

On that site i quoted it says that the property is -(a+b)=(-a)+(-b) wher a and b are 2 numbers so we could fit say 5 amd 7 for a and b giving -(5 + 7) = (-5) + (-7) -12 = -12

OpenStudy (welshfella):

- it look like D to me what do you think?

OpenStudy (welshfella):

in D a = -6 and b = 17

OpenStudy (anonymous):

So that property is when an equation does something like this? (3+5) - (3) + (5)